OpenStudy (anonymous):

if z is 4(cos80+ i sin80) .. how to find i/z?

OpenStudy (mertsj):

Write i in the form 0+1i and write it in its polar form

OpenStudy (mertsj):

0+i=1(cos90+isin90)\[\frac{1(\cos 90+i \sin 90)}{4(\cos 80+i \sin 80)}=\frac{1}{4}(\cos (90-80)+i \sin (90-80))=\]

OpenStudy (mertsj):

\[\frac{1}{4}(\cos 10+i \sin 10)\]

OpenStudy (anonymous):

ok thanks...

OpenStudy (mertsj):

yw